solve by elimination
2x-3y=1
12x-18y=6
show work

Answer by Prithwis(166)

2x-3y=1 .....(multiply with -6)
=> -12x + 18y = -6
12x-18y = 6
-------------------
Add: 0 = 0
The system of equations ar dependent and so, all real values of x and y satisfy the equations.